Find out if the inverse exists for the following, give reasoning behind your answer. If you conclude that the inverse exists then find the B ́ezout coefficients and the inverse of the modulo. [Hint: Example 2 of section 4.4 in the book]


(a) - (3 points) 678 modulo 2970

(b) - (3 points) 137 modulo 2350


1
Expert's answer
2022-03-08T10:12:02-0500

Solution


According to, Euclidean Algorithm, the inverse of a mod b exists, if and only if

"g.c.d {\color{Red} (a, b)} = 1"


Here g.c.d stands for the greatest common divisor.

Solution (a)


To check the inverse for 678 modulo 2970 exists or not, we can write


"2978 = 678 × 4 + 258\\
678 = 258 × 2 + 162\\
258 = 162 × 1 + 96\\
162 = 96 × 1 = 66\\
96 = 66 × 1 + 30\\
66 = 30 × 2 + 6 \\
30 = 6 × 5 + 0 \\"


Hence "g.c.d (2978, 678) = 6 \u2260 1"


Therefore, the inverse of 678 modulo 2970 does not exist.  



Solution (a)


To check the inverse for 137 modulo 2350 exists or not, we can write


"2350 = 137 × 17 + 21\\

137 = 21 × 6 + 11\\

21 = 11 × 1 + 10\\

11 = 10 × 1 + 1 \\"


Hence "g.c.d (2350, 137) = 1"


Therefore, an inverse of 137 modulo 2350 exists.  


Now to find the inverse we proceed as follows.


"11 + 10 (– 1) = 1\\

21 + 11 (– 1) = 10\\

137 + 21 (– 6) = 11\\

2350 + 137 (– 17) = 21"


Therefore,

"11 + (21 + 11 (– 1)) (– 1) = 1\\

11 – 21 + 11 = 1 \\

11 × 2 + 21 (-1) = 1"


Then

"(137 + 21 (-6)) × 2 + 21 (-1) = 1\\

137 × 2 + 21 × (-13) = 1"

 

And finally

"137 × 2 + (2350 + 137 × (-17))(-13) =1 \\

2350 × 2(-13) + 137 × 2(223) = 1"


Hence Bezout coefficients are (-13) and (223)

 

And the invers of 137 modulo 2350 is "223"